A professor wants to supplement her pension with investment interest. If she invests $31,000 at 6% interest, how much more would she have to invest at 7% to achieve a goal of $2,560 per year in supplemental income?

Answer by mananth(16946)

31000 at 6%
so interest = 0.06 * 31000
=$1860
Total interest to be earned = $2560
Balance to be earned = $2560-$1860
=$700
..
I= p*r*t
700=p*0.07*1
700/0.07 =p
p= $10,000 This is the investment required.